OpenStudy (loser66):

Old problem confirm \[\sum_{n=1}^\infty \dfrac{z^{n^2}}{n}=\sum_{k=1}^\infty a_kz^k\] Then \(a_k =?? \) Please, help

OpenStudy (loser66):

@SithsAndGiggles

OpenStudy (anonymous):

If you're replacing \(n^2\) with \(k\), that means you have this pattern: \[\begin{array}{c|cccc} n&1&2&3&4&5&\cdots\\ \hline k&1&4&9&16&25&\cdots \end{array}\] So while \(a_n\) is defined for all positive integers \(n\), \(a_k\) is only defined for perfect squares. In other words, you can transform \(a_n\) to \[a_k=\begin{cases}a_{\sqrt n}&\text{ for }n\in\{1,4,9,16,25,\ldots\}\\0&\text{otherwise}\end{cases}\]

OpenStudy (loser66):



OpenStudy (loser66):